Find the missing coordinate to complete the ordered pair so that the point ( , 1) is on the line 2x + y = - 15.

Answer by alicealc(293)

y = 1
2x + y = -15
2x + 1 = -15
2x = -15 - 1
2x = -16
x = -16/2 = -8

so, the missing coordinate is x = -8